OpenStudy (anonymous):

Let X and Y be natural numbers. Prove that x^2+y+2 and y^2 +4x cannot be both perfect squares. Please help

OpenStudy (anonymous):

I do not have an idea to start this

OpenStudy (amistre64):

hmmm, it says nothing about the same perfect square number does it

OpenStudy (anonymous):

I guess not

OpenStudy (amistre64):

what class is this for?

OpenStudy (anonymous):

This is a maths Olympiad question (1990 I think)

OpenStudy (amistre64):

lets assume they are and prove that its not

OpenStudy (amistre64):

is the sum of perfect squares a perfect square? what property do perfect squares have that we can demonstarte?

OpenStudy (amistre64):

9+4 = 13 so the sum isnt

OpenStudy (anonymous):

I tried to do somthing with a^2-b^2= (a-b)(a+b) but got nowhere

OpenStudy (amistre64):

a^2 + b^2 = c^2 is the sum of perfect squares as a perfect square right?

OpenStudy (anonymous):

Nope, that is Fermat's little theorem. Right?

OpenStudy (anonymous):

Or that is with cubes? :) Was long time ago

OpenStudy (amistre64):

i was thinking pythag thrm, but yeah.

OpenStudy (amistre64):

fermats is a^n + b^n = c^n for n>2

OpenStudy (anonymous):

Oh yeah :) Never mind, that is offtopic now

OpenStudy (amistre64):

start with a if p then q statement

OpenStudy (amistre64):

see if the contrapositive is easier to determine, maybe

OpenStudy (amistre64):

let x^2+y+2 = a^2 let y^2+4x = b^2 x^2+y+2 + y^2+4x = c^2 (x^2+4x)+ (y^2+y) = c^2 -2 (x^2+4x+4)+ (y^2+y+1/4) = c^2 -2+4+1/4 (x+2)^2+ (y+1/2)^2 = c^2 + 2.25 dunno if this gets us anywhere, but its about the only thought i got at the moment

OpenStudy (anonymous):

x^2+y+2 + y^2+4x = c^2 This is not necessarily true. Only for Pythagoras triplets

OpenStudy (amistre64):

i agree, but can we define x and y so that it make a pythag triple?

OpenStudy (anonymous):

That is not given, so I guess we cannot.

OpenStudy (amistre64):

we just need to prove it false, so one counter example is all we need

OpenStudy (anonymous):

No because it says that we only need one good X, Y (anything)

OpenStudy (amistre64):

then for a start, think about properties of perfect squares then apply that property to what we have to see if it holds

OpenStudy (amistre64):

pythag has perfect squares so i thought it might be prudent. im open to being wrong of course

OpenStudy (rational):

i remember working this problem the other day in mse

OpenStudy (anonymous):

Well thanks for trying, I have to go now. Will think on it tomorrow. It bothers me though

OpenStudy (amistre64):

i dont remember working it at all in over 40 years :)

OpenStudy (rational):

hold up, it is easy

OpenStudy (amistre64):

maybe: a^2/b^2 = (a/b)^2 ?

OpenStudy (anonymous):

I found the problem: http://www.imomath.com/index.php?options=Hi&mod=23&ttn=Hungary-Israel No solutions there...

OpenStudy (anonymous):

Smart

OpenStudy (amistre64):

some property of perfect squares :) just needed to determine the property

OpenStudy (amistre64):

does it matter than not all k=0,1 (mod4) are perfect squares?

OpenStudy (amistre64):

ill trust you on that for now lol

OpenStudy (rational):

@amistre64 i see my stupid mistake, it wont work yeah :/ will get back to this after dinner..

OpenStudy (amistre64):

see, i trust you to find your way ;)

OpenStudy (rational):

:) found this http://artofproblemsolving.com/community/c6h392528_prove_there_are_no